wiki:UC20200901

Version 1 (modified by trac, 6 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 09/01/2020

Attending: Nancy Chanover (NMSU), Jamey Eriksen (APO), Bill Ketzeback, John Bally (CU), Moire Prescott (NMSU), Eric Bellm (UW), Kevin Schlaufman (JHU), Anne Verbiscer (UVa), Ben Williams (UW), Misty Bentz (GSU), Denise Stephens (BYU), Russet McMillan (APO), Mukremin Kilic (OU), Joanne Hughes (Seattle U), Jeff Bary (NAPG), Sarah Tuttle (UW), Kal Kadlec (UW)


User feedback and comments from institutional representatives:

  • Wyoming - no report
  • Washington - nothing to report
  • UVa - nothing to report
  • Seattle - nothing to report
  • Oklahoma - nothing to report
  • NMSU - nothing to report
  • NAPG - nothing to report
  • JHU - nothing to report
  • FGCU - nothing to report
  • Georgia State - nothing to report
  • Colorado - nothing to report
  • BYU - nothing to report

Telescope and Instruments Report:

Jamey's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 7/30/2020 - 8/27/2020

0) Overview

APO has been open for night operations since June 1. We are still restricting all visitors to the site. Following the state of New Mexico and NMSU guidelines, there are travel restrictions for staff as well as visitors. Currently, as NMSU is in Phase 3 of the Return to Research, APO is allowed between 35-50% of staffing. The NMSU numbers, coupled with the state of New Mexico restrictions for a non-essential classified business like APO, which allow staffing levels of 50% of the Fire Marshall determined maximum occupancy of a building, has allowed us to have more people on site at any given time. However, those restrictions are still limiting work activities for day time engineering tasks. Weather for the month of August has been as expected for monsoon rain season. We are losing time consistently since the beginning of the month due to weather. We are experiencing an above average year of Miller moths and as such are still dealing with them. Happily, the moth issue has abated considerably since the heavier monsoon rains have started. Keeping fingers crossed that the issue doesn’t get worse again in the fall.

1) Telescope

The telescope has been performing as expected. The previously noted issue with the tertiary rotation (post shutdown) has not been an issue during August, as such, this issue won’t be reported on going forward.

The telescope exhaust fan system was repaired/put back into operations and hasn’t had a reoccurrence of the issue.

2) Instruments

DIS: Scattered light is still an issue with the instrument. Turret/grating has been showing improved behavior lately. It has been mostly working on the first/second attempt at movement. It will be monitored to catch if it gets worse.

TripleSpec: powered on, cold and operational.

Agile: powered on, cold and operational.

Echelle: powered on, cold and operational. The ratio used (in-order light/inter-order light) to track scattered light is currently stable for blue and red.

NIC-FPS: Frequent expose inits are being required to remedy an all-zero readout problem. The corrupted readouts are not new but seem to have changed their pattern, apparently coming in clusters sometimes and not at other times. Previous known triggers such as instrument handling or rotating the instrument upside down may not be as likely to trigger the issue now, but we haven't really figured out what - if anything - is a common trigger. Investigations are ongoing.

ARCTIC: powered on, cold, and operational. Issue with dome flats is being debugged and solutions are being worked on.

3) ARCSAT ARCSAT is operational and has not been used for much observations this past month. Weather is the main limiting factor. A low flux level streaking issue in SurveyCam images is being investigated.


Additional telescope and instrument discussion:

No updates regarding telescope or instruments beyond what was included in the monthly report. The site is still operating at a reduced on-site staffing level due to COVID restrictions. Jamey described the protocols in the event that a member of the site staff tests positive for COVID-19. In summary, the observatory would shut down for a week or two, during which time we would conduct extensive contact tracing to determine who was at risk for exposure and also do a deep cleaning.


ARCTIC Flats

We are still working to mitigate the problem of dome flats not being very flat with the new NA2 baffle in place. Tonight during engineering time the observing staff will experiment with some changes to the baffle. In parallel with that effort we are still working on building up a library of good sky flats. Joanne explained that for her science, which requires very accurate photometry, it is not feasible to use twilight flats if the sky conditions are partly cloudy. For her upcoming run she will be using the Washington C filter and then standard SDSS and BVRI filters; she will try to plan ahead and communicate her filter choices early so that we can attempt to populate the sky flat library accordingly. Russet indicated that she would like to do super flats, especially in the redder filters where fringing is an issue, but those are processed images and thus perhaps not appropriate for every kind of investigation. Joanne asked if a script exists that could enable one to dither and get the flats that way. Russet will investigate. John Bally reported having good success with the extended dither script. Although it is intended for NICFPS, the guider is the same as for ARCTIC, so in principle it should work. Russet will test to see if it works for certain offset types. Kevin asked if the flat structure is a simple 2D gradient that depends on the orientation of the instrument? Russet replied that it is more complicated: it goes across most of the field and then it flattens out, so it is not a linear gradient. As expected, the gradient gets steeper when the instrument sees more light from the mirror covers. In most rotations, the gradient is seen across most of the chip and then it flattens out. We still recommend that users try to use sky flats whenever possible.


APO and the COVID-19 virus

The New Mexico Governor's public health order still mandates a 14-day self-quarantine for anyone traveling into New Mexico from out of state. Her latest orders lifted some restrictions on in-restaurant dining. Things may ease up further in mid-September, depending on whether the case numbers continue to decrease after Labor Day weekend. Of relevance to our users, the main issues are (a) we are still not allowing visitors to the site due to the 14-day quarantine, and (b) we are still operating at a reduced staffing level and with additional cleaning protocols, so please continue to exercise patience with the observing staff.


Q3 3.5m scheduling:

All time in Q3 has been assigned.


Q3 ARCSAT scheduling:

There is some OPEN ARCSAT time (4.8 nights on 21-25 September) that users can request. It would have to be with SurveyCam, using the specific filters that were in place during the preceding program.

UC reps: have your users let Ben, Nancy and Russet know if they are interested in requesting any of the remaining ARCSAT time.


Q4 3.5m scheduling:

Russet is actively working on building the Q4 schedule. As expected, October A halves are oversubscribed. There are some scheduling conflicts that are still being resolved.

Nancy discussed the fact that the quarterly institutional passwords, which are mailed out from APO via USPS before the start of each quarter, have not been reaching some people since they are working from home and not checking (or cannot access) their work mailboxes. Nancy requested that anyone who would like their institutional password mailed to their home address should send the info to Nancy; she will compile them and pass the addresses on to the APO staff.


Updated Training Policies and Procedures:

As discussed last month, we are going to offer remote training sessions for individuals at ARC member institutions or leasing partners who need to be trained for remote observing on the 3.5m. This does not replace the class trainings, which should be requested via the standard proposal process. Rather, it is intended to address the one-off cases where an individual needs to be trained. Russet has tentatively penciled in a sequence of B halves in mid-Oct around Oct 10 for the Q4 training session, but that is still subject to change as the Q4 schedule is still being built. We currently have 10-11 people signed up, which is about the maximum we can handle for a single group. The BYU training will be scheduled separately. We are eager to get feedback on how this goes since we intend to offer this opportunity on a quarterly basis as long as travel restrictions remain in place.


KOSMOS Update:

Sarah expects to receive an update on the dewar status later today. After that she will be developing a delivery plan that either includes the dewar being fixed or consists of putting the detector back in the old dewar so the instrument can be shipped to APO for commissioning. The TEG still has limited access to their lab at UW. Misty asked that the tentative schedule for KOSMOS commissioning be distributed to the Users Committee as soon as it is known, since there are programs that depend on the long-slit capability.


ARCTIC Diffuser Rotation

The TEG has received the parts it ordered and can run the rotating mechanism for the diffuser in the lab after preliminary assembly. The parts were sent out for anodization, and will be reassembled when they come back. At that point they will do environmental testing and then ship it to APO, hopefully by then end of September.


DAWG Update

Nancy expects to receive the DAWG white paper early this month, and hopes to be able to report on it at next month's meeting.


Science Communications Update

Nancy and Emily are working to plan another Science Hangout with a modified format based on the feedback they received from this group last month. In the mean time, they would like to post some short turnaround snippets on social media that describe how the telescope was used on a given night. They will contact the program PIs ahead of time to ask for permission and solicit any sample images, spectra, guider images, screenshots of the observing team, or anything else that the users wish to provide. The PIs may also elect NOT to participate in this, or can dictate the amount of information that is shared with the public (e.g. no target names). Nancy and Emily also welcome unsolicited submissions for this effort! UC reps: please give your users a heads-up that they may see some of these requests in the near future.


Planetary Decadal Survey White Paper

Nancy will be submitting a white paper for the Planetary Science and Astrobiology Decadal Survey 2023-2032 that argues for the continued relevance of 4m-class telescopes for planetary science investigations. The link is here if 3.5m users wish to co-sign or provide suggested edits to the document: https://docs.google.com/document/d/1TySsjeqvHYw-Gq6rcgMe-FpwskDBlVJEyhBtXctGbD8/edit?usp=sharing. All edits must be received by Monday September 14 as white papers are due on September 15.


ACTION ITEMS:

Open action items from previous meetings:

  • ARCTIC filter order - Nancy. Status: OPEN. Still awaiting some pricing info from vendors.
  • UC reps: notify users that there is a limited amount of OPEN/DD time remaining on both the 3.5m and ARCSAT during Q3. If interested in requesting some time, please follow standard protocols for making such requests. Status: CLOSED for 3.5m, but OPEN for ARCSAT.
  • UC reps: remind all users to use twilight flats with ARCTIC rather than dome flats when possible, and to consider contributing flats to the APO library of ARCTIC flats that is being built up (contact Russet McMillan for details). Status: CLOSED.
  • UC reps: Remind users that Nancy needs science highlights from each institution for the ARC Board of Governors meeting. They are due to Nancy no later than October 15, and can be collected by a single representative from each institution and mailed as a packet (easier for Nancy!) or sent to her directly by individual investigators. There is no slide template, but each slide should contain at minimum: project title, author(s) and affiliation(s), a graphic, a few text bullets describing the result, and a citation if there is one. Status: OPEN.

New action items from this meeting:

  • UC reps: send Nancy an alternate address if you want your quarterly institutional passwords to be mailed somewhere other than your workplace.
  • UC reps: alert your users to the fact that they may be asked for contributions to post on social media.
  • UC reps: advertise the planetary science decadal survey white paper; the deadline for contributions or signatories is Sep 14.
  • Nancy: resend the meeting schedule for Fall 2020 (see All Other Business below).

All Other Business

Clarifications on the Users Committee meeting schedule for Fall 2020:

a) The meetings will all be on Tuesdays.

b) The November and December meetings will be on the SECOND Tuesday of the month.

Thus, the remaining meetings in 2020 will be on October 6, November 10, and December 8.


Next meeting:

The next meeting will be on Tuesday October 6 at 10:30 am MDT.


Note: See TracWiki for help on using the wiki.